OpenStudy (anonymous):

A cylinder’s radius is reduced to 2/5 its original size and the height is quadrupled. How has the volume of the cylinder changed? A. The volume is now 8/25 the original volume. B. The volume is now 16/25 the original volume. C. The volume is now 64/25 the original volume. D. The volume is now 16/5 the original volume.

OpenStudy (campbell_st):

well if you let radius = 1 and height = h the volume of the cylinder is \[V = \pi \times1^1\times h = \pi h\] so if the radius becomes 2/5 and height becomes 4h then \[V = \pi \times (\frac{2}{5})^2 \times 4h\] or \[V=\frac{16}{25} \pi h\] what happened to the volume...